What is the cubic feet of a cylinder that measures a radius of 11" by a height of 13"? It is a washing machine drum.
Answer by nyc_function(2741)

V = pi*r^2*h
V = (3.14)(11)^2 * 13
Just multiply across to find V or volume of cylinder.
=====================================================
V = pi*r^2*h
V = (3.14)(11)^2 * 13

In other words, you want me to do it for you. How is this going to help you when facing similar questions alone in the classroom?
V = 4939.22 inches.
There are 12 inches in a foot. So, there are 12^3 inches^3 (cubic inches) in a foot^3 (cubic foot). Multiply things out using inches, then divide by 12^3.
If this does not work for you, search theinternet for a free converter website that converts inches to cubic feet.